OpenStudy (anonymous):

break-even - how many do I have to sell to break even if, 55/15-4x=break-even point?? $55=fixed cost, $15=selling price, and $4x=variable costs Please help!! I'm stumped on trying to figure this out??

OpenStudy (anonymous):

Someone Please Help Me?? Please??

OpenStudy (anonymous):

I solved my own problem Thank You!! Here it is!! 55+4x=15x because, 55+4x-4=15x-4, and 55/11=11x/11, so, x=5 . YAY! I DID IT!!